OpenStudy (helpsos):

evaluate: -7^0 (-7)^0

OpenStudy (anonymous):

Anything to the power of 0 is 1.

OpenStudy (helpsos):

up for the first one my teacher marked me wrong since i wrote one

OpenStudy (anonymous):

\(-7^0\) is \(-1\) because subtraction comes after powers.

OpenStudy (helpsos):

i wrote -7^0 = 1 but she marked it wrong

OpenStudy (anonymous):

\[ -7^0 =-(7^0) = -(1) = -1 \]

OpenStudy (helpsos):

oh ok

OpenStudy (helpsos):

tahnx :)

OpenStudy (anonymous):

Sometimes it gets tricky.

OpenStudy (helpsos):

is the answer for (-7)^0 also -1?

OpenStudy (anonymous):

No, it's just 1

OpenStudy (helpsos):

ok thanx and yeah this is tricky for me :P i h8

OpenStudy (helpsos):

but thanx again :D